Am explorer walks 22km due North and then walks 45degrees south east for 47km.what’s the resultant displacement from the origin?


1
Expert's answer
2022-03-16T17:01:33-0400

Suppose explorer start from origin "A" to north "22Km" to a point "B".

Then Move "45\degree" South East "47Km" to reach a point "C".

Join point "AC," Length of "AC" would be displacement from origin




Now draw a line "DC" from point "C" that meet line "AB" on point "D"

"\triangle ABC" Given

"\angle B= 45 \degree" "AB=22Km"

"BC=47 Km"


"Displacement =AC=?"



From "\triangle DBC"

"SinB=\frac{DC}{BC}"

"Sin45 \degree=\frac{DC}{47}"

"DC={BC}{Sin45 \degree}=47Sin45 \degree"

"DC=33.2340Km"


"\triangle DBC"

"TanB=\frac{DC}{BD}"

"Tan45 \degree=\frac{DC}{BD}"

"BD=DC=33.23Km"


"AD=BD-AB=33.23-22"

"AD=11.23Km"

"\triangle ADC"

"{AC}^2={DC}^2 +{AD}^2"

"{AC}^2={33.23}^2 +{11.23}^2"

"{AC}^2=1230.3458"

"AC=\sqrt{1230.3458}"

"AC=35.0049 Km"

"Displacement\boxed {AC=35.00 Km}"